### v1

@@ V-B Comparison Between Indoor and Outdoor Humidity Mixing Ratios
The previous analysis has shown that RH is excessively high in the basement and Room 5 on the ground floor. The high humidity in the basement is mainly due to the earthen floor, which allows groundwater continuously evaporates from the wet soil into the indoor air. This moisture is speculated to rise through the walls, acting as a moisture source that affects not only the basement but also rooms on the upper floors. As depicted in Fig. 10a, MRs in the basement and ground floor rooms were consistently higher than outdoors, suggesting an internal moisture source. The problem of high humidity is particularly pronounced in Room 5, where the sensor box is positioned on the floor and near the walls. The rising dampness has led to high humidity in Room 5. Thus, the evaporation from the basement need to be handled to take control of this problem. One measure is to install vapor barriers in the basement to reduce evaporation and moisture migration.
@@ V-D Short-term Fluctuation of Relative Humidity
Since rooms on the first floor maintain relatively stable RH levels, continuous heating throughout the winter might be unnecessary. Implementing more energy-efficient heating strategies, such as intermittent heating, can ensure human comfort during occupancy—especially during long gatherings in Room 103 with large groups—while also conserving energy.
@@ V-E Impact of Precipitation on Groundwater Level Changes
The earthen floor of the basement contains abundant groundwater, which is evidently influenced by precipitation. Fig. 14 illustrates the effect of hourly outdoor precipitation on the hourly changes in GWL at two monitoring points in Room 014 of the basement. Channel 2, located on the east side, and Channel 1, on the west side, showed different responses to rainfall.
@@ V-F Impact of Occupants on Indoor Climate
Löfstad Castle is open for public activities. The presence of occupants can impact indoor climate. Fig. 16 illustrates changes in five indoor environmental parameters—temperature, RH, CO2 concentration, dust concentration, and noise level—in Room 103 over nine days from December 2 to 10, 2023. Visitor gatherings occurred during the daytime on December 2, 3, 9, and 10. During these gatherings, all five environmental parameters increased considerably. The highest CO2 concentration on these days even reached or exceeded 3000 ppm, which could cause discomfort for occupants [65]. Besides, there seems to have maintenance work on December 7, as indicated by a significant rise in dust concentration and a slight increase in noise levels (see Fig. 16d and Fig. 16e).
@@ VI Conclusion
This study contributes to the field of historic building conservation by demonstrating a comprehensive digitalization solution and its application. The solution and findings not only apply to Löfstad Castle but also offer important insights and practical guidance for the conservation of other historic buildings facing similar problems.

### v2

@@ V-B Comparison Between Indoor and Outdoor Humidity Mixing Ratios
The previous analysis has shown that RH is excessively high in the basement and Room 5 on the ground floor. The high humidity in the basement is mainly due to the direct soil flooring, which allows groundwater continuously evaporates from the wet soil into the indoor air. This moisture is speculated to rise through the walls, acting as a moisture source that affects not only the basement but also rooms on the upper floors. As depicted in Fig. 11a, MRs in the basement and ground floor rooms were consistently higher than outdoors, suggesting an internal moisture source. The problem of high humidity is particularly pronounced in Room 5, where the sensor box is positioned on the floor and near the walls. The rising dampness has led to high humidity in Room 5. Thus, the evaporation from the basement need to be handled to take control of this problem. One measure is to install vapor barriers in the basement to reduce evaporation and moisture migration.
@@ V-D Short-term Fluctuation of Relative Humidity
Since rooms on the first floor maintain relatively stable RH levels, continuous heating throughout the winter might be unnecessary. Implementing more energy-efficient heating strategies, such as intermittent heating, can ensure human comfort during occupancy—especially during long gatherings in Room 103 with large groups—while also conserving energy. Additionally, a zoned heating approach could allow for targeted heating in different areas of the building based on conservation needs and occupancy patterns. This strategy not only conserves energy but also improves occupant comfort by providing customized climate control in frequently used spaces while maintaining lower temperatures in less occupied areas.
@@ V-E Impact of Precipitation on Groundwater Level Changes
In the basement, the exposed soil of the floor contains abundant groundwater, which is evidently influenced by precipitation. Fig. 15 illustrates the effect of hourly outdoor precipitation on the hourly changes in GWL at two monitoring points in Room 014 of the basement. Channel 2, located on the east side, and Channel 1, on the west side, showed different responses to rainfall.
@@ V-F Impact of Occupants on Indoor Climate
Löfstad Castle is open for public activities. The presence of occupants can impact indoor climate. Fig. 17 illustrates changes in five indoor environmental parameters—temperature, RH, CO2 concentration, dust concentration, and noise level—in Room 103 over nine days from December 2 to 10, 2023. Visitor gatherings occurred during the daytime on December 2, 3, 9, and 10. During these gatherings, all five environmental parameters increased considerably. Pearson correlation analysis confirmed the impact of occupant presence on environmental parameters. As shown in Fig. 18, CO2 concentration had a strong positive correlation with temperature ($r=0.6$) and dust concentration ($r=0.6$), and a very strong positive correlation with RH ($r=0.81$) and noise level ($r=0.83$). The highest CO2 concentration on these days even reached or exceeded 3000 ppm, which could cause discomfort for occupants [67]. Besides, there seems to have maintenance work on December 7, as indicated by a significant rise in dust concentration and a slight increase in noise levels (see Fig. 17d and Fig. 17e).
@@
Figure 18: The matrix of Pearson correlation coefficients between five indoor environmental parameters of Room 103 from December 2 to 11, 2023.
@@ VI Conclusion
This study contributes to the field of historic building conservation by demonstrating a comprehensive digitalization solution and its application. The scalable and adaptable solution supports diverse historic buildings beyond Löfstad Castle. Its modular design allows the integration of additional sensors and data sources, making it suitable for diverse architectural layouts and conservation needs. Using cloud-based storage and analysis, the system can handle large datasets from complex buildings, enabling data-driven conservation. Furthermore, adaptability of the solution allows customization for specific environmental monitoring needs and heritage contexts. It provides conservation practitioners with applicable tools for tracking environmental conditions, analyzing long-term trends, and implementing sustainable preservation measures. The solution and findings from this study thus offer insights and practical guidance for the conservation of other historic buildings facing similar challenges.